\begin{thebibliography}{99}

\bibitem{am1} M. H. Ahmadi Gandomani and M. J. Mehdipour, Generalized derivations on some convolution algebras. Aequationes Math. 92 (2018), no. 2, 223--241.

\bibitem{am} M. H. Ahmadi Gandomani and M. J. Mehdipour, Symmetric bi-derivations and their generalizations on group algebras. Filomat 35 (2021), no. 4, 1233--1240.

\bibitem{dl} H. G. Dales and A. T. Lau, The second duals of Beurling algebras, Mem. Amer. Math. Soc., 177 (2005), no. 836, 1--191. 

\bibitem{d} Q. Deng,  On a conjecture of Vukman. Internat. J. Math. Math. Sci. 20 (1997), no. 2, 263--266.

\bibitem{lp} A. T. Lau and J. Pym, Concerning the second dual of the group algebra of a locally compact group. J. London Math. Soc. (2) 41 (1990), no. 3, 445--460.

\bibitem{mnr}  S. Maghsoudi, R. Nasr-Isfahani and A. Rejali,  Arens multiplication on Banach algebras related to locally compact semigroups. Math. Nachr. 281 (2008), no. 10, 1495--1510.

\bibitem{m1} G. Maksa,  A remark on symmetric biadditive functions having nonnegative diagonalization. Glasnik Mat. Ser. III 15(35) (1980), no. 2, 279--282.

\bibitem{m2} G. Maksa, On the trace of symmetric bi-derivations. C. R. Math. Rep. Acad. Sci. Canada 9 (1987), no. 6, 303--307.

\bibitem{mm}  M. J. Mehdipour and Gh. R. Moghimi,, The existence of nonzero compact right multipliers and Arens regularity of weighted Banach algebras. Rocky Mountain J. Math. 52 (2022), no. 6, 2101--2112.

\bibitem{mms} M. J. Mehdipour, Gh. R. Moghimi and N. Salkhordeh, Jordan derivations on certain Banach algebras, arXiv: 2306.12529.

\bibitem{mn} M. J. Mehdipour and R. Nasr-Isfahani, Completely continuous elements of Banach algebras related to locally compact groups. Bull. Austral. Math. Soc. 76 (2007), no. 1, 49--54.

\bibitem{ms} M. J. Mehdipour and Z. Saeedi,  Derivations on convolution algebras. Bull. Korean Math. Soc. 52 (2015), no. 4, 1123--1132.

\bibitem{p} C. Park,  Biderivations and bihomomorphisms in Banach algebras. Filomat 33 (2019), no. 8, 2317--2328.

\bibitem{r0} H. Reiter and J. D. Stegeman, Classical Harmonic Analysis and Locally Compact Groups, Second edition, London Mathematical Society Monographs, New Series, 22, The Clarendon Press, Oxford University Press, New York, 2000.

\bibitem{v1} J. Vukman, Symmetric bi-derivations on prime and semi-prime rings. Aequationes Math. 38 (1989), no. 2-3, 245--254.

\bibitem{v2} J. Vukman, Two results concerning symmetric bi-derivations on prime rings. Aequationes Math. 40 (1990), no. 2-3, 181--189.

\bibitem{w} Y. Wang, Biderivations of triangular rings. Linear Multilinear Algebra 64 (2016), no. 10, 1952--1959. 
\end{thebibliography}

\documentclass[a4paper,12pt,reqno]{article}
\usepackage[T1]{fontenc}
\usepackage{authblk}
\usepackage{epsfig,graphicx,subfigure}
\usepackage{amsthm,amsmath,amssymb,amsfonts}
\usepackage{color,xcolor}
\usepackage[all]{xy}
\usepackage[top=40mm, bottom=40mm, left=40mm, right=35mm]{geometry}
\usepackage{hyperref}

\newtheorem{theorem}{Theorem}[section]

\newtheorem{corollary}[theorem]{Corollary}
\theoremstyle{definition}

\newtheorem{example}[theorem]{Example}

\newtheorem{remark}[theorem]{Remark}
\numberwithin{equation}{section}

\begin{document}
\setcounter{page}{1}
\title{\vspace{-1.5cm}
\vspace{.5cm}
\vspace{.7cm}
{\large{\bf  Symmetric bi-derivations on certain Banach algebras} }}
 \date{}
\author{{\small \vspace{-2mm}  M. Eisaei$^1$ and Gh. R. Moghimi $^1$\footnote{Corresponding author} }}
\affil{\small{\vspace{-4mm}  $^1$ Department of Mathematics,
Payame Noor University (PNU), Tehran 19395-4697, Iran }}
\affil{\small{\vspace{-4mm}  mojdehessaei59@student.pnu.ac.ir }}
\affil{\small{\vspace{-4mm} moghimimath@pnu.ac.ir}}
\maketitle
\hrule
\begin{abstract}
\noindent
Let $A$ be a Banach algebra with a right identity $u$ such that $uA$ is commutative and semisimple. In this paper, we investigate symmetric bi-derivations of $A$ and detremine their range. We also study symmetric bi-derivations of $A$ with their $k$-centralizing trace. Finally, we prove every symmetric Jordan bi-derivation of $A$ is a symmetric bi-derivation.
 \end{abstract}

 \noindent \textbf{Keywords}: Bi-derivation, $k$-centralizing, $k$-skew centralizing.\\
{\textbf{2020 MSC}}:  47B47, 16W25.
\\
\hrule
\vspace{0.5 cm}
\baselineskip=.8cm
\section{Introduction}

Let $A$ be a Banach algebra.
A mapping $D: A \times A \rightarrow A$ is called a \textit{symmetric bi-linear}  mapping  if $D(a,b)=D(b,a)$,
\begin{equation*}
D( \alpha a, b)= \alpha D(a,b) \quad \text{and} \quad D(a+b,c)=D(a,c)+ D(b,c).
\end{equation*}
for all $a,b,c \in A$ and $\alpha \in \Bbb{C}$. Also,
$D$ is called a \textit{symmetric Jordan bi-derivation} if
$D$ is symmetric bi-linear and
\begin{equation*}
D(a^2,b)= D(a,b) a + a D(a,b)
\end{equation*}
for all $a,b \in A$.
Furthermore, if for every $a,b,c \in A$, we have
\begin{equation*}
D(ab, c)= D(a,c) b + a  D(b,c), 
\end{equation*}
then $D$ is called a \textit{symmetric bi-derivation}.
Clearly, every Jordan bi-derivation is a bi-derivation.
But, the converse is not true, in general.
The mapping $f: A \rightarrow A$ defined by
\begin{equation*}
f(a)=D(a,a)
\end{equation*}
is called the \textit{trace of } $D$.

Let $T: A \rightarrow A$ be a mapping and $k \in \Bbb{N}$. Then $T$ is called 
\textit{$k$-centralizing \emph{(}resp. $k$-skew centralizing\emph{)}}  if for every $a \in A$,
\begin{equation*}
[T(a), a^k]:= T(a) a^k - a^k T(a) \,\,\, (\text{resp. } \langle T(a), a^k \rangle= T(a) a^k + a^k T(a) )
\end{equation*}
is an element of $Z(A)$, the set of all $a\in A$ such that $ax=xa$ for all $x\in A$. In particular, if for every $a \in A$
\begin{equation*}
[T(a), a^k]=0, \,\,\, (\text{resp.} \langle T(a), a^k \rangle =0).
\end{equation*}
then $T$ is called \emph{$k$-commuting} (resp. \emph{$k$-skew commuting}). 

Maksa \cite{m1, m2} introduced and studied symmetric bi-derivations.
Some authors continued this investigation \cite{am, d, p, v1, v2, w}.
For example, vukman \cite{v2} showed that if
$D: R \times R \rightarrow R$ is a symmetric bi-derivation on a non commutative prime ring of characteristic note two and there $R$ such that for every $r \in R$,
\begin{equation*}
[[f(r),r],r] \in Z(R),
\end{equation*}
then $D$ is zero.

In this paper, we always assume that $A$ is a Banach algebra with a right identity $u$ such that $uA$ is commutative and semisimple. We prove that every symmetric bi-derivation of $A$ maps $A$ into $\hbox{ran}(A)$. We also show that if the trace of a symmetric bi-derivation of $A$ is $k$-centralizing, then the symmetric bi-derivation is zero. Finally, we establish that every symmetric Jordan bi-derivation of $A$ is a symmetric bi-derivation.
\section{Main Results}
 Before we give the first our result, let us recall that the right annihilator of $A$ is denoted by $\hbox{ran}(A)$ and it is the set of all $r\in A$ such that $ar=0$ for all $a\in A$. Note that $\hbox{ran}(A)$ is a subset of the radical of $A$.

\begin{theorem} \label{m1}
Let $D: A \times A \rightarrow A$ be a symmetric bi-derivation. Then the range of $D$ is contained into $ \emph{ran}(A)$.
\end{theorem}
\begin{proof}
First, note that  the range of $D$ is equal to
$\cup_{m \in A} \phi_m(A)$,
where $$\phi_m(n)= D(m,n)$$ for all
$n \in A$.
But $\phi_m$ is a derivation on $A$ and so
$\phi_m(A) \subseteq  \hbox{ran}(A)$; see \cite{mms}.
This proves the theorem.
\end{proof}
\begin{remark}
Let $\mathcal{A}$ be a Banach algebra such that $\frac{\mathcal{A}}{\hbox{ran}(A)}$ is commutative and semisimple.
Then a similar argument to the proof of Theorem 2.1 shows that the range of any symmetric bi-derivation $D: \mathcal{A} \times \mathcal{A} \rightarrow \mathcal{A}$ is contained into
$ \hbox{ran}(\mathcal{A})$.
\end{remark}
\begin{theorem} \label{m2}
Let $D: A \times A \rightarrow A$ be a symmetric bi-derivation and $f$ be the trace of $D$.
Then the following assertions are equivalent.

\emph{(a)} There exists $k \in \Bbb{N}$ such that $f$ is $k$-commuting.

\emph{(b)} There exists $k \in \Bbb{N}$ such that $f$ is $k$-centralizing.

\emph{(c)} There exists $k \in \Bbb{N}$ such that $f$ is $k$-skew commuting.

\emph{(d)} There exists $k \in \Bbb{N}$ such that $f$ is $k$-skew centralizing.

\emph{(e)} $D=0$.
\end{theorem}
\begin{proof}
Let $m \in A$ and $k \in \Bbb{N}$.
It follows from Theorem 2.1 that
\begin{eqnarray*}
[D(m,m), m^k]&=& D(m,m) m^k - m^k D(m,m) \\
&=& D(m,m) m^k + m^k D(m,m) \\
&=& \langle D(m,m), m^k \rangle.
\end{eqnarray*}
Thus
\begin{equation*}
[f(m), m^k] = \langle f(m), m^k \rangle.
\end{equation*}
So the statements (a)-(d) are equivalent.
Assume now that $f$ is $k$-commuting.
Then
\begin{eqnarray} \label{po1}
f(m) m^k &=& D(m,m) m^k  \nonumber \\
&=& D(m,m) m^k - m^k D(m,m) \nonumber  \\
&=& [D(m,m), m^k]\\
&=& [f(m), m^k]\nonumber\\
&=&0\nonumber
\end{eqnarray}
for all $m \in A$.
Hence $f(u)=0$. We also have
\begin{eqnarray*}
f(m+u)&=& D(m+u, m+u) \nonumber \\
&=& D(m,m) +2 D(m,u)+ D(u,u) \nonumber  \\
&=& f(m) + 2 D(m,u)
\end{eqnarray*}
for all $m \in A$. This together with (2.1) shows that
\begin{eqnarray*}
0&=& f(m+u) (m+u)^k \\
&=& (f(m)+ 2 D(m,u)) (m+u)^k \\
&=& (f(m)+ 2 D(m,u)) \bigg ( \sum_{j=0}^{k-1} \binom{k-1}{j}m^{k-j} + \sum_{j=0}^{n-1} \binom{n-1}{j} u m^{k-j-1} \bigg)
\end{eqnarray*}
for all $m \in A$. This implies that
\begin{eqnarray} \label{po3}
 \sum_{j=1}^{k} \binom{k}{j} f(m) m^{k-j} +2 \sum_{j=0}^{k} \binom{k}{j} D(m,u) m^{k-j}  =0.
\end{eqnarray}
Put
\begin{eqnarray*}
E_1(m)=  \sum_{ \tiny{\begin{array}{c}
j \, even  \\
 j=2
\end{array}}}^{k} \binom{k}{j} f(m) m^{k-j}, \quad
E_2(m)= 2  \sum_{ \tiny{\begin{array}{c}
j \, even  \\
 j=0
\end{array}}}^{k} \binom{k}{j} D(m,u) m^{k-j},
\end{eqnarray*}
\begin{eqnarray*}
O_1(m)=  \sum_{ \tiny{\begin{array}{c}
j \, odd  \\
 j=1
\end{array}}}^{k} \binom{k}{j} f(m) m^{k-j}, \quad
O_2(m)= 2  \sum_{ \tiny{\begin{array}{c}
j \, odd  \\
 j=1
\end{array}}} \binom{k}{j} D(m,u) m^{k-j}.
\end{eqnarray*}
It follows from (2.2) that
\begin{equation} \label{op4}
E_1(m) + E_2(m) + O_1(m)+ O_2(m)=0
\end{equation}
for all $m \in A$.
If we replace $m$ by $-m$ in (2.3), then
\begin{equation*}
E_1(m) - E_2(m) - O_1(m)+ O_2(m)=0
\end{equation*}
Hence
\begin{equation} \label{op5}
E_1(m)+O_2(m)=0
\end{equation}
and so
\begin{equation} \label{op6}
O_1(m)+ E_2(m)=0.
\end{equation}
Now, let $k$ be even. Then by (2.5), we have
\begin{eqnarray*}
0 &=& O_1(m) + E_2(m) \\
&=&   \sum_{ \tiny{\begin{array}{c}
j \, odd  \\
 j=1
\end{array}}}^{k-1} \binom{k}{j} f(m) m^{k-j} +
2 \sum_{ \tiny{\begin{array}{c}
j \, even   \\
 j=0
\end{array}}}^{k-2}
\binom{k}{j}  D(m,u) m^{k-j} + 2 D(m,u).
\end{eqnarray*}
Thus for every $r \in \hbox{ran} (A)$, we obtain
\begin{equation} \label{op7}
D(r,u)=0,
\end{equation}
because $$f(r) r^{k-j} = D(r,u) r^{k-j}=0.$$
Since $m-um \in \hbox{ran}(A)$, it follows that
\begin{eqnarray*}
0&=& D(m-um,u) \\
&=& D(m,u) - D(u,u) m- uD(m,u) \\
&=& D(m,u) - f(u) m \\
&=& D(m,u)
\end{eqnarray*}
for all $m \in A$.
Hence $O_2=0$ and thus
\begin{equation} \label{op8}
\sum_{ \tiny{\begin{array}{c}
j \, even   \\
 j=2
\end{array}}}^{k}
\binom{k}{j}  f(m) m^{k-j}=0.
\end{equation}
If $i$ is even and $2 \leq i \leq k-2$, then by
(2.7), we have
\begin{eqnarray*}
0&=&
\sum_{ \tiny{\begin{array}{c}
j \, even   \\
 j=2
\end{array}}}^{k}
\binom{k}{j}  f(m) m^{k+i-j} \\
&=& \sum_{ \tiny{\begin{array}{c}
j \, even   \\
 j=2
\end{array}}}^{k}
\binom{k}{j}  f(m) m^{k} m^{i-j} +
 \sum_{ \tiny{\begin{array}{c}
j \, even   \\
 j=i+2
\end{array}}}^{k}
\binom{k}{j}  f(m) m^{k+i-j}\\
&=& \sum_{ \tiny{\begin{array}{c}
j \, even   \\
 j=i+2
\end{array}}}^{k}
\binom{k}{j}  f(m) m^{k+i-j}.
\end{eqnarray*}
For $i=k-2$, we get
$f(m) m^{k-2} =0$.
Hence
\begin{equation*}
\sum_{ \tiny{\begin{array}{c}
j \, even   \\
 j=4
\end{array}}}^{k}
\binom{k}{j}  f(m) m^{k-j}=0\quad\hbox{ and} \quad
\sum_{ \tiny{\begin{array}{c}
j \, even   \\
 j=i+2
\end{array}}}^{k}
\binom{k}{j}  f(m) m^{k+i-j}=0.
\end{equation*}
Continuing this procedure, we have $f=0$ and therefore $D=0$.
Similarly, if  $k$ is odd, one can prove that $D=0$.
\end{proof}
\begin{theorem} \label{m4}
Let $D: A \times A \rightarrow A$ be a symmetric Jordan bi-derivation.
Then $D$ is a symmetric bi-derivation.
\end{theorem}
\begin{proof}
Let $m \in A$ and $\phi_m$ be the function defined as the proof of Theorem 2.1. Then for every $n \in A$, we have
\begin{eqnarray*}
\phi_m(n^2) &=& D(m,n^2) \\
&=& D(m,n) n + n D(m,n)\\
& =& \phi_m(n) n + n \phi_m(n).
\end{eqnarray*}
Thus $\phi_m$ is a Jordan derivation on $A$ for all $m \in A$.
Hence, $\phi_m$ is a derivation on $A$, see Theorem 2.2 of \cite{mms}. Therefore,
\begin{eqnarray*}
D(m k, n) &=& \phi_n(mk) \\
&=& \phi_n(m) k + m \phi_n (k) \\
&=& D(m,n)k + m D(k,n)
\end{eqnarray*}
for all $m,n,k \in A$.
That is, $D$ is a symmetric bi-derivation.
\end{proof}

As an immediate consequence of Theorems 2.1 and 2.4 we have the following result.

\begin{corollary} Let ${\cal A}$ be a unital commutative semisimple Banach algebra. Then the zero map is the only symmetric bi-derivation on ${\cal A}$.
\end{corollary}

We finsh the paper with the following example.

\begin{example} Let $\omega$ be a weight function on $[0, \infty)$. 

(i) Let $M(\omega)$ be the Banach algebra of all complex regular Borel measures $ \mu $ on $[0, \infty)$ such that $ \omega \mu  \in M( [0, \infty)) $, the Banach algebra of all complex regular Borel measures on $ [0, \infty)$. It is well-known that $M(\omega)$ is a unital commutative semisimple Banach algebra; see \cite{dl, r0}. Hence  the only symmetric bi-derivation on $M(\omega)$ is zero.

(ii) Let $L_0^\infty(\omega)$ be the Banach space of all Lebesgue measure functions $f$ on $[0, \infty)$ such that
\begin{equation*}
\lim_{x \rightarrow \infty} \text{ess sup}
 \bigg \{ \frac{f(y) \chi_{(x, \infty)} (y)}{w(y)}: y \geq 0 \bigg\}=0,
\end{equation*}
where $\chi_{(x, \infty)}$ is the characteristic function of $(x, \infty)$ on $[0, \infty)$. Then $L_0^\infty(\omega)^*$ is a Banach algebra \cite{mnr}. Also, $L_0^\infty(\omega)^*$ has a right identity, say $u$, and $uL_0^\infty(\omega)^*$ is isometrically isomorphic to $M(\omega)$; for an extensive study of this Banach algebra see \cite{am1, lp, mnr, mm, mn}. One can prove that the radical of $L_0^\infty(\omega)^*$ is equal to $\hbox{ran}(L_0^\infty(\omega)^*)$; see \cite{ms}. Therefore, every symmetric Jordan bi-derivation of $L_0^\infty(\omega)^*$ is a symmetric bi-derivation and its range is into the radical of $L_0^\infty(\omega)^*$. Also, there is no non-zero symmetric bi-derivation with $k$-cerntralizing trace.
\end{example}

\end{document}